# On Selection of Cross-Section Averages in Non-stationary Environments

### v2

@@ 6.3 Part $M\subset M_{0}$
$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
@@
$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\left[(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right]T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
@@
$\leq\underbrace{N^{-1}T^{-1}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\sqrt{N}\mathbf{F}_{\mathbf{x}}\overline{\mathbf{V}}\right\|^{2}\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|^{2}}_{O_{p}((NT)^{-1})}$
@@
$+\underbrace{2T^{(\tau-1)/2}N^{-1/2}\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\sqrt{N}\mathbf{F}_{\mathbf{x}}\overline{\mathbf{V}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|^{2}\left\|\overline{\boldsymbol{\Gamma}}\right\|}_{O_{p}(T^{(\tau-1)/2}N^{-1/2})}$
@@
$+\underbrace{T^{\tau}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right\|\left\|T^{-(1+\tau)}\mathbf{F}_{x}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|^{2}\left\|\mathbf{q}_{M}\right\|^{2}\left\|\overline{\boldsymbol{\Gamma}}\right\|^{2}\frac{1}{N}\sum_{i=1}^{N}\left\|\boldsymbol{\Gamma}_{i}\right\|^{2}}_{O_{p}(T^{\tau}\xi_{N,T,\tau}^{-1})}$
$=O_{p}(T^{\tau}\xi_{N,T,\tau}^{-1}),$ (A.18)
@@
which is a slowly decaying term dominated by $T^{(\tau-1)/2}=o(1)$, since $\tau\in(0,1)$. We let $C_{N,T}=\mathrm{min}(\sqrt{N},\sqrt{T})$, and we move on to
@@
$\left\|\mathbf{I}_{c}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
@@
$=\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
@@
$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
@@
$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
@@
$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\left[(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right]T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
@@
$\leq\underbrace{T^{-(1+\tau)/2}N^{-1/2}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|\left\|T^{-(1+\tau)/2}\sqrt{N}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|}_{o_{p}(C_{N,T}^{-2})}$
@@
$+\underbrace{\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|\overline{\boldsymbol{\Gamma}}\right\|}_{O_{p}(C_{N,T}^{-2})}$
@@
$+\underbrace{N^{-1/2}T^{-1}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|\sqrt{N}T^{-(1+\tau)/2}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\overline{\boldsymbol{\Gamma}}\right\|\left\|\boldsymbol{\Gamma}_{i}\right\|}_{O_{p}(N^{-1/2}T^{-1})}$
@@
$+\underbrace{T^{(\tau-1)/2}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|T^{-(1+\tau)/2}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|}_{T^{(\tau-1)/2}\xi_{N,T,\tau}^{-1})}$
@@
×$\times\underbrace{\left\|\overline{\boldsymbol{\Gamma}}\right\|^{2}\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|}_{O_{p}(1)}$
$=O_{p}(C_{N,T}^{-2}),$ (A.19)
@@
$\left\|\mathbf{I}_{d}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{V}_{i}\right\|$
@@
$\leq\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{V}_{i}\right\|$
@@
$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|$
@@
$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{V}_{i}\right\|$
@@
$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\left[(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right]T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|$

### v3

@@ 5.3.1 Case of $M\subset M_{0}$
$+\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
@@
$+\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\left[(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right]T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
@@
$\leq\underbrace{N^{-1}T^{-1-\tau}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\sqrt{N}\mathbf{F}_{\mathbf{x}}\overline{\mathbf{V}}\right\|^{2}\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|^{2}}_{O_{p}(N^{-1}T^{-1-\tau})}$
@@
$+\underbrace{2T^{-(\tau+1)/2}N^{-1/2}\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\sqrt{N}\mathbf{F}_{\mathbf{x}}^{\prime}\overline{\mathbf{V}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|^{2}\left\|\overline{\boldsymbol{\Gamma}}\right\|}_{O_{p}(T^{-(\tau+1)/2}N^{-1/2})}$
@@
$+\underbrace{\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right\|\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|^{2}\left\|\mathbf{q}_{M}\right\|^{2}\left\|\overline{\boldsymbol{\Gamma}}\right\|^{2}\frac{1}{N}\sum_{i=1}^{N}\left\|\boldsymbol{\Gamma}_{i}\right\|^{2}}_{O_{p}(\xi_{N,T,\tau}^{-1})}$
$=O_{p}(\xi_{N,T,\tau}^{-1})=o_{p}(T^{-\tau}),$ (A.19)
@@
which remains a slowly decaying term even if scaled by $T^{\tau}$, because it will be dominated by $T^{(\tau-1)/2}=o(1)$ stemming from the second term, since $\tau\in(0,1)$. We let $C_{N,T}=\mathrm{min}(\sqrt{N},\sqrt{T})$, and we move on to
@@
$\left\|\mathbf{I}_{c}\right\|$ $=\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
@@
$=\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
@@
$+\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
@@
$+\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
@@
$+\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\left[(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right]T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
@@
$\leq\underbrace{T^{-(1+3\tau)/2}N^{-1/2}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|\left\|T^{-(1+\tau)/2}\sqrt{N}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|}_{O_{p}(T^{-(1+3\tau)/2}N^{-1/2}C_{N,T}^{-2})}$
@@
$+\underbrace{T^{-\tau}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|\overline{\boldsymbol{\Gamma}}\right\|}_{O_{p}(T^{-\tau}C_{N,T}^{-2})}$
@@
$+\underbrace{N^{-1/2}T^{-(1+\tau)}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|\sqrt{N}T^{-(1+\tau)/2}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|\mathbf{q}_{M}\right\|^{2}\left\|\overline{\boldsymbol{\Gamma}}\right\|\left\|\boldsymbol{\Gamma}_{i}\right\|}_{O_{p}(N^{-1/2}T^{-1-\tau})}$
@@
$+\underbrace{T^{-(\tau+1)/2}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\right\|\left\|T^{-(1+\tau)/2}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|}_{O_{p}(T^{-(\tau+1)/2}\xi_{N,T,\tau}^{-1})}$
@@
$\times\underbrace{\left\|\overline{\boldsymbol{\Gamma}}\right\|^{2}\left\|\mathbf{q}_{M}\right\|^{2}\left\|\boldsymbol{\Gamma}_{i}\right\|}_{O_{p}(1)}$
$=O_{p}(T^{-\tau}C_{N,T}^{-2})+O_{p}(N^{-1/2}T^{-1-\tau})=o_{p}(T^{-\tau}),$ (A.20)
@@
$\left\|\mathbf{I}_{d}\right\|$ $=\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{V}_{i}\right\|$
@@
$\leq\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{V}_{i}\right\|$
@@
$+\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|$
@@
$+\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{V}_{i}\right\|$
@@
$+\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\left[(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right]T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|$